# Enhanced indexation using both equity assets and index options

### v1

@@ 3.5 Option strategies and optimisation
So consider the situation where we make use of a standard enhanced indexation optimisation approach, but include option strategies (artificial assets) evaluated in-sample using past return behaviour as discussed above.
It can arise that for some option strategies we never actively trade (buy/sell) an option when that strategy is evaluated in-sample. But, as illustrated by the example in Table 2, any investment allocated to that strategy will increase (e.g. if $u>1$) due to the risk-free rate. Logically this means that such an option strategy has no contribution to make in-sample (beyond accumulating cash at the risk-free rate).
@@
Optimising using in-sample data to decide the proportion of our wealth to be allocated to each asset (equity and option strategy) we may find that no investment is made in some option strategies (artificial assets). That is, given past history, these strategies do not seem to be worthwhile. However there may be one (or more) option strategies that are allocated some proportion of our wealth. Equally there may also be some equity assets that are allocated some proportion of our wealth. So based on in-sample data we have what would have been a good past investment using both equity assets and option strategies.
So now what do we do for the (unknown) future? For any equity asset we simply buy the asset (using the proportion of wealth allocated to it) and hold it unchanged into the future.
For any option strategy we (using the proportion of wealth allocated to it) apply it into the future. However here the situation with an option strategy is more complicated than for equity assets as the option strategy may (depending upon how it is constructed and future market conditions) never actually buy any options in the future. So the wealth allocated to it is left untouched and never used (although the wealth allocated to it may increase at the risk-free rate, as in the example in Table 2). But this is perfectly acceptable, if the option strategy was a good one to apply in the past, and we have now applied it as the unknown future revealed itself, the fact that it never buys any options is perfectly acceptable (after all, better not to buy than buy and lose wealth).
Obviously there are no guarantees here, what was a good equity asset to hold in the past may not be a good equity asset to hold into the future. The same applies to option strategies.
3.6 Contribution and extensions
Above we introduced the concept of an option strategy, which enables us to include options as an artificial asset in enhanced indexation. We believe that the concept of an option strategy, not presented previously in the academic literature as best as we are aware, is the key intellectual contribution of this paper. As will become apparent in the computational results presented later below introducing option strategies in an enhanced indexation setting offers clear benefits in terms of improved out-of-sample performance.
However we believe that the idea of an option strategy is not just restricted to S&P 500 index options, such as considered in this paper, but has a significant number of extensions:
  • we can introduce equity asset call and put options, e.g. those that are traded on major companies represented in the index
  • we can consider other types of options, e.g. futures
  • we can consider any portfolio optimisation problem that involves equity assets, e.g. Markowitz (efficient frontier) approaches
  • we can produce approaches that only trade options over time in an attempt to make money
Since an option strategy, as presented in this paper, is simply an artificial asset some of the above extensions are obvious, namely introducing equity asset call and put options; considering other types of options and Markowitz (efficient frontier) approaches. To use option strategies to produce approaches that only trade options over time in an attempt to make money we merely restrict the asset universe to just contain option strategies.
@@ 3.7 Choice of an optimisation based enhanced indexation approach
We would strike a note of caution here that, because one might consider many different option strategies, there is a danger of in-sample over-fitting. By this we mean suppose we generate millions of different option strategies. When we use optimisation to decide which ones to use on historic data (so in-sample) the optimiser will naturally attempt to maximise its objective, choosing from the millions of strategies those that make best use of in-sample data.
However when we invest in these option strategies, i.e. apply them out-of-sample into the future, we may find that by over-fitting in-sample we do not get good out-of-sample performance. In many respects this is not an issue when we use just equity assets, since the number of such assets is typically limited by the number of equity assets in the index. In the results presented below we restricted ourselves to just a few option strategies to avoid any dangers of over-fitting.
@@ 4 Second-order stochastic dominance
For brevity here we assume that the reader has some familiarity with SSD. Readers unfamiliar with this topic are referred to Fabian et al. (2011a, b); Roman et al. (2006).
@@
  • $r_{it}$ be the return of asset $i\in E\cup F\cup S$ at time $t$, where for assets in $E\cup F$ these returns are computed from market data, for the option strategies (artificial assets) in $S$ these returns are computed by applying each strategy as discussed above
@@
In Equation (9) $\kappa_{s}=1$ for unscaled SSD and $\kappa_{s}=(s/T)$ for scaled SSD. Equation (7), in conjunction with Equation (9), maximises the minimum (scaled) tail difference. Scaling, due to Fabian et al. (2011b), gives more importance to the returns in the right tails of the distribution. Since in this paper we are seeking to outperform the index emphasising the returns in the right-tail seems appropriate, so we only use scaled SSD in our computational work below.
@@ 5 Computational experiments
In this section we first discuss the asset dataset used, which is associated with equities in the S&P 500 index. We then discuss the option strategies we generated, and finally we show computational results illustrating the benefits of employing option strategies in an enhanced indexation setting. In our computational results we consider investment in S&P 500 equities and investment in a S&P 500 exchange-traded fund both with, and without, option strategies.
@@ 5.1 Dataset
The main dataset consists of daily stock prices from 17th March 2017 until 1st August 2025. This time period, over 8 years, includes the Covid pandemic, which had a significant effect on stock prices. Our data has been manually adjusted to account for survivorship bias - on a given date only assets that were part of the S&P 500 index at that time are available to be selected for investment. In order to define the scenarios required by SSD we used a lookback approach that included the most recent 201 daily prices, which then yield 200 in-sample returns.
@@
In this paper we deal with S&P 500 options. We would anticipate that option prices as quoted in the market might differ from their theoretical BS price. However market participants have told us that for liquid (commonly traded) options the BS price is a good approximation of the market quoted price. In addition we lack access to historical data as to daily option prices over our time period (2017-2025). For these reasons we, in this paper, use BS to price the options we consider.
Finally, we used the time series for the VIX index (https://finance.yahoo.com/quote/%5EVIX/) as an approximation for implied-volatility for use in BS option pricing. We used the IRX (https://finance.yahoo.com/quote/%5EIRX) which represents the yield of the US 13-week Treasury bill, as a proxy for short-term risk-free investment and for pricing options with BS. The original IRX data is annualised, for use here we produce daily returns using the formula $r_{\text{daily}}=(1+r^{\frac{1}{252}})-1$.
@@ 5.2 Option strategies
As an addition to the asset universe described above (S&P 500 equities and SPY), we generated 12 option strategies based on three different policies. They are the following:
@@
    Buy/keep a {straddle, strangle} when the annualised realised S&P 500 volatility over the past {20, 30} days is below 8%. Straddles and strangles are non-directional trading strategies designed to profit from sudden increases in volatility. A straddle buys an equal position in ATM calls and puts, while a strangle buys an equal position in OTM calls and puts. As for moneyness we used a 3% OTM target. These four option strategies bet on a “storm after a calm period”.
For all of these strategies the rollover rule was to rollover if option lifetime satisfies $L<20$ days or if moneyness has deviated significantly from the $3\%$ OTM target, so $100|0.97F_{t}-E|/0.97F_{t}\geq 3$. Note here that with a rollover rule of $L<20$ days we never hold any option to expiry.
The economic logic behind the first two of these strategies is:
@@
The economic logic behind the last of these three strategies would entail too much explanation to be worthwhile here. We would merely state that the use of straddles and strangles is well recognised in option trading.
@@ 5.3 Investment in equities
For numeric insight into these strategies, we show in Figure 1 a chart displaying their cumulative out-of-sample performance and in Table 3 some selected comparative statistics. These are calculated from the out-of-sample returns of the two strategies, and correspondingly for the S&P 500 index.
It is clear from Figure 1 that SSD with options significantly outperforms SSD without options.
@@
Let $Q$ be a series of $0,\ldots,T$ daily portfolio values, where $Q_{t}$ is the value of the given portfolio on day $t$. In Table 3 FV stands for the final portfolio value, assuming a starting amount of \$1, and is calculated as $Q_{T}/Q_{0}$. CAGR stands for Capital Annualised Growth Rate and as a percentage is calculated as $100\left(\left(\frac{Q_{T}}{Q_{0}}\right)^{\frac{1}{Y}}-1\right)$, where $Y=T/252$ is an approximation for the number of years in the out-of-sample period. Sharpe and Sortino are the annualised Sharpe and Sortino ratios respectively, where for their calculation we use the CBOE 10-year treasury notes (symbol TNX) as the risk-free rate. For these four performance measures high values are better.
@@ 5.4 Investment in SPY
Passive investing through the purchase of ETFs and index funds has been increasing. In 1993, passive funds invested in US stocks managed \$23 billion of assets, 3.7% of the combined assets managed by active and passive funds, and 0.44% of the US stock market. By 2021, passive assets had risen to \$8.4 trillion, 53% of combined active and passive and 16% of the stock market Jiang et al. (2025).
In this section we evaluate whether it would be profitable to adopt a semi-passive strategy which uses the scaled SSD model with an asset universe composed of just the SPY ETF and the 12 option strategies. Figure 3 shows out-of-sample cumulative returns, and Table 4 displays selected out-of-sample statistics. Both follow the same format as Figure 1 and Table 3.
Again we see that using (scaled) SSD with options performs best.
@@
Detailed examination of the results revealed that during that particular period of high option exposure in 2022, the in-sample selected option strategies typically did not “activate” out-of-sample, i.e. there were few out-of-sample option trades meaning most out-of-sample investment was just in the risk-free asset, hence the non-volatile performance during a short period in Figure 3. As noted previously above if necessary exposure limits can easily be added to limit option exposure via use of Equation (14).
@@ 7 Conclusions
In this paper we considered how we can include index options in enhanced indexation. We presented the concept of an option strategy which for a known set of options is a specified set of rules which detail how these options are to be traded depending upon market conditions. An option strategy overcomes the fundamental difficulty that individual options only have a short life and enables us to treat an option strategy as an artificial asset.
We demonstrated the use of option strategies in enhanced indexation by using an enhanced indexation approach based on second-order stochastic dominance. We considered index options for the S&P 500, using a dataset of daily stock prices from 2017-2025 that has been manually adjusted to account for survivorship bias. This dataset has been made publicly available for use by future researchers.

### v2

@@ 3.5 Option strategies and optimisation
So consider the situation where we make use of a standard enhanced indexation optimisation approach, but include option strategies evaluated in-sample using past return behaviour as discussed above.
It can arise that for some option strategies we never actively trade (buy/sell) an option when that strategy is evaluated in-sample. But, as illustrated by the example in Table 2, any investment allocated to that strategy will increase, e.g. if $u>1$, due to the risk-free rate. Logically this means that such an option strategy has no contribution to make in-sample, beyond accumulating cash at the risk-free rate.
@@
Optimising using in-sample data to decide the proportion of our wealth to be allocated to each asset (equity and option strategy) we may find that no investment is made in some option strategies. That is, given past history, these strategies do not seem to be worthwhile. However there may be one, or more, option strategies that are allocated some proportion of our wealth. Equally there may also be some equity assets that are allocated some proportion of our wealth. So based on in-sample data we have what would have been a good past investment using both equity assets and option strategies.
So now what do we do for the unknown future? For any equity asset we simply buy the asset, using the proportion of wealth allocated to it, and hold it unchanged into the future.
For any option strategy we, using the proportion of wealth allocated to it, apply it into the future. However here the situation with an option strategy is more complicated than for equity assets as the option strategy may, depending upon how it is constructed and future market conditions, never actually buy any options in the future. So the wealth allocated to it is left untouched. Note here however that the wealth allocated to it may increase at the risk-free rate, as in the example in Table 2.
@@ 3.6 Choice of an optimisation based enhanced indexation approach
We would strike a note of caution here that, because one might consider many different option strategies, there is a danger of in-sample over-fitting. By this we mean suppose that we produce millions of different option strategies. When we optimise to decide which ones to use based on historic data (so in-sample) the optimiser will naturally attempt to maximise its objective, choosing from the millions of strategies those that make best use of in-sample data.
However when we invest in these option strategies, i.e. apply them out-of-sample into the future, we may find that by over-fitting in-sample we do not get good out-of-sample performance. In many respects this is not an issue when we just use equity assets, since the number of such assets is typically limited by the number of equity assets in the index. In the results presented below we restricted ourselves to just a few option strategies to avoid any dangers of over-fitting.
@@ 4 Second-order stochastic dominance
For brevity here we assume that the reader has some familiarity with SSD. Readers unfamiliar with this topic are referred to Fábián et al. (2011a, b); Roman et al. (2006).
@@
  • $r_{it}$ be the return of asset $i\in E\cup F\cup S$ at time $t$, where for assets in $E\cup F$ these returns are computed from market data, for the option strategies in $S$ these returns are computed by applying each strategy as discussed above
@@
In Equation (9) $\kappa_{s}=1$ for unscaled SSD and $\kappa_{s}=(s/T)$ for scaled SSD. Equation (7), in conjunction with Equation (9), maximises the minimum (scaled) tail difference. Scaling, due to Fábián et al. (2011b), gives more importance to the returns in the right tail of the distribution. Since in this paper we are seeking to outperform the index emphasising the returns in the right tail seems appropriate, so we only use scaled SSD in our computational work below.
@@ 5 Computational experiments
In this section we first discuss the asset dataset used, which is associated with equities in the S&P 500 index. We then discuss the option strategies we used, and show computational results illustrating the benefits of employing option strategies in an enhanced indexation setting. We go on the discuss performance in different market regimes (bull/bear) and the potential impact of transaction cost. In our computational results we consider investment in S&P 500 equities and investment in a S&P 500 exchange-traded fund both with, and without, option strategies.
@@ 5.1 Dataset
The main dataset consists of daily stock prices from 17th March 2017 until 1st August 2025. This time period, over 8 years, includes the Covid pandemic, which had a significant effect on stock prices. Our data has been manually adjusted to account for index composition. So on a given date only assets that were part of the S&P 500 index at that time are available to be selected for investment, so we do not make use of any knowledge as to assets that might in the future become part of the S&P 500 (or cease to be in the S&P 500). In order to define the scenarios required by SSD we used a lookback approach that included the most recent 201 daily prices, which then yield 200 in-sample returns.
@@
In this paper we deal with S&P 500 options. We would anticipate that option prices as quoted in the market might differ from their theoretical BS price. However for liquid (commonly traded) options the BS price is “widely used” (Brenner, 2025). In addition we lack access to historical data as to daily option prices over our time period (2017-2025). For these reasons we, in this paper, use BS to price the options we consider.
Finally, we used the time series for the VIX index (https://finance.yahoo.com/quote/%5EVIX/) as an approximation for implied volatility for use in BS option pricing. We used the IRX (https://finance.yahoo.com/quote/%5EIRX) which represents the yield of the US 13-week Treasury bill, as a proxy for short-term risk-free investment and for pricing options with BS. The original IRX data is annualised, for use here we produce daily returns using the formula $r_{\text{daily}}=(1+r^{\frac{1}{252}})-1$, where 252 represents the number of trading days in a year.
@@ 5.2 Option strategies
As an addition to the asset universe described above (S&P 500 equities and SPY), we made use of 12 option strategies based on three different policies. They are the following:
@@
    Buy/keep a {straddle, strangle} when the annualised realised S&P 500 volatility over the past {20, 30} days is below 8%. These four option strategies bet on a “storm after a calm period”.
For all of these strategies the rollover rule was to rollover if option lifetime satisfies $L<20$ days or if moneyness has deviated significantly from the $3\%$ OTM target, so $100|0.97F_{t}-E|/0.97F_{t}\geq 3$. Note especially here that with this lifetime ($L$) rollover rule we never hold any option to expiry.
The economic logic behind these strategies is:
@@
  3. 3.
    Straddles and strangles are non-directional trading strategies designed to profit from increases in volatility (Hull, 2021). A straddle buys an equal position in ATM calls and puts, with identical lifetimes and strike prices. It achieves a positive payoff when the underlying asset experiences sufficiently large price movements in either direction. A strangle is similar, with the main difference being the purchase of a call and a put with different strike prices, typically with both options being OTM.
@@ 5.3 Investment in equities
For numeric insight into these strategies, we show in Figure 1 a chart displaying their cumulative out-of-sample performance and in Table 3 some selected comparative statistics. These are calculated from the out-of-sample returns of the two strategies, and correspondingly for the S&P 500 index. It is clear from Figure 1 that SSD with options significantly outperforms SSD without options.
@@
Let $Q$ be a series of $0,\ldots,T$ daily portfolio values, where $Q_{t}$ is the value of the given portfolio on day $t$. In Table 3 FV stands for the final portfolio value, assuming a starting amount of \$1, and is calculated as $Q_{T}/Q_{0}$. CAGR stands for Capital Annualised Growth Rate and as a percentage is calculated as $100\left(\left(\frac{Q_{T}}{Q_{0}}\right)^{\frac{1}{Y}}-1\right)$, where $Y=T/252$ is an approximation for the number of years in the period. Sharpe and Sortino are the annualised Sharpe and Sortino ratios respectively, where for their calculation we use the CBOE 10-year treasury notes (symbol TNX) as the risk-free rate. For these four performance measures high values are better.
@@
5.4 Performance in different regimes
In this section, we evaluate the out-of-sample performance of both strategies and the S&P 500 conditional on market regimes. We consider both bullish (prices rising) and bearish (prices falling) markets. We manually classified the S&P 500 price history as either bull or bear according to the visually observed price trend. Figure 3 includes four panels and shows both the visual regime classifications and the out-of-sample performance across each regime.
(a) Periods classified as bull market
(b) Out-of-sample performance during bull periods
(c) Periods classified as bear market
(d) Out-of-sample performance during bear periods
Figure 3: Out-of-sample performance across different regimes
In panels (a) and (c), we highlight (using vertical shading) the periods that are manually classified as either bull or bear markets; each period is assigned to exactly one regime. Panels (b) and (d) present regime-conditional out-of-sample performance. We constructed each regime-specific return series by concatenating the strategies and index returns corresponding to the highlighted days into a set of continuous asset return time series. Note here that panel (b) involves over 1400 days, panel (d) only around 350 days.
Regime Strategy FV CAGR Sharpe Sortino CVaR Vol MDD
Bull market SSD: equities with options 5.29 31.11 1.32 1.97 2.81 20.09 16.87
SSD: equities without options 4.78 28.97 1.27 1.82 2.82 19.42 16.87
S&P 500 5.34 31.32 1.74 2.62 2.08 14.79 9.60
Bear market SSD: equities with options 0.68 -23.68 -0.81 -1.10 4.45 31.10 39.99
SSD: equities without options 0.60 -30.06 -1.15 -1.55 4.50 29.92 46.91
S&P 500 0.43 -44.70 -1.65 -2.16 5.15 34.12 60.75
Table 4: Out-of-sample statistics across different market regimes
Comparative regime-conditional statistics are shown in Table 4. Unsurprisingly, during periods classified as bullish, the S&P 500 showed strong performance and relatively low risk. The two SSD strategies had similar risk profiles, higher than the index, but the strategy with options achieved better returns, almost matching that of the S&P 500. During periods classified as bearish, both strategies strongly outperformed the S&P 500, both in terms of risk and return. SSD with options out-performed SSD without options on all but one measure.
Overall, the addition of option strategies enhanced the performance of SSD both by increasing upside potential during “good” times and mitigating downside risk during “bad” times.
@@ 5.5 Investment in SPY
Passive investing through the purchase of ETFs and index funds has been increasing. In 1993 passive funds invested in US stocks managed \$23 billion of assets, some 0.44% of the US stock market. By 2021 passive assets had risen to \$8.4 trillion, some 16% of the stock market (Jiang et al., 2025).
In this section we evaluate whether it would be profitable to adopt a semi-passive strategy which uses the scaled SSD model with an asset universe composed of just the SPY ETF and the 12 option strategies. Figure 4 shows out-of-sample cumulative returns, and Table 5 displays selected out-of-sample statistics. Both follow the same format as Figure 1 and Table 3. Again we see that using (scaled) SSD with options performs best.
@@
Detailed examination of the results revealed that during the particular period of high option exposure in 2022, the in-sample selected option strategies typically did not “activate” out-of-sample, i.e. there were few out-of-sample option trades meaning most out-of-sample investment was just in the risk-free asset, hence the non-volatile performance during a short period in Figure 4. As noted previously above if necessary exposure limits can easily be added to limit option exposure via use of Equation (14).
@@
5.6 Estimated transaction costs
[…]
Let $f$ be the average fractional turnover over the 91 rebalances in our work. Then if $\alpha$ is the fractional loss in value on the portfolio traded at each rebalance the value of $\alpha$ that reduces the final value of the portfolio to that of the S&P 500 at the end of the period is given by
$(1-\alpha f)^{91}\text{(FV of our approach)}=\text{(FV of the S\&P~500)}$ (15)
So for example for the first case in Table 3, SSD with equities and options, the average proportion of the portfolio traded was $f=0.4881$, and we have a FV for our approach of 3.61 and a FV for the S&P 500 of 2.31. From Equation (15) $\alpha=0.0100$, equivalently 100 basis points (bps). So if average transaction cost per rebalance is less than this the FV from our approach will exceed the FV of the S&P 500. Whilst recognising that transaction costs vary according to timing and size of trade this value of 100 bps exceeds values commonly quoted for institutional trading (15-30 bps for a one-way equity trade). Hence it seems likely from this ex-post analysis that even in the presence of transaction costs our approach would result in a final value exceeding that of the S&P 500.
For SSD without options in Table 3 $f=0.4856$, giving $\alpha=0.0051$ (51 bps), whilst for SSD with SPY and options in Table 5 $f=0.0794$ giving $\alpha=0.0247$ (247 bps), based on the FV for SPY of 2.61.
Whilst the above analysis neglects any transaction cost resulting from option strategy trading in each out-of-sample period between rebalances detailed examination of the results indicated that this would not significantly alter the values for $\alpha$ seen above. For example as can be seen visually in Figure 2 investment in option strategies only occasionally exceeds 10% of the portfolio, and is often substantially less. With respect to Figure 5, where the investment in options exceeds that seen in Figure 2, note that the $\alpha$ value is much higher, 247 bps as compared with 100 bps.
@@ 7 Conclusions
In this paper we considered how we can include index options in enhanced indexation. We presented the concept of an option strategy which for a given option is a specified set of rules which detail how the option is to be traded depending upon market conditions. An option strategy overcomes the fundamental difficulty that individual options only have a short life and enables us to treat an option strategy as equivalent to an asset.
We demonstrated the use of option strategies in enhanced indexation by using an enhanced indexation approach based on second-order stochastic dominance. We considered index options for the S&P 500, using a dataset of daily stock prices from 2017-2025 that has been manually adjusted to account for index composition. This dataset has been made publicly available for use by future researchers.